A baggage handler drops your 7.90 kg suitcase onto a conveyor belt running at 1.50 m/s . The materials are such that μs = 0.570 and μk = 0.250. How far is your suitcase dragged before it is riding smoothly on the belt?

How far is your suitcase dragged before it is riding smoothly on the belt?
Answer #1

Sum of the horiz. forces = m*a = -friction
a = -friction/m
Sum of the vert. forces = 0 = N - m*g
N = m*g

friction = u*N = u*m*g
a = -uk*m*g/m = -uk*g

I am assuming that the suitcase is moving so kinetic friction will be used. Once the bag has stopped moving, then you can use static friction,
a = -uk*g
a = -0.25*9.81 = -2.452 m/s^2

using vf^2 = vi^2 + 2*a*d
vf = 0
d = -vi^2/(2*a) = -(1.50^2)/(2*-2.45) = 0.445 m